OpenStudy (anonymous):

What is the maximum amount of Li2CO3 that can form when 3.13 mol of CO2 reacts with excess LiOH? CO2+2LiOH yields Li2CO3+H2O A. 1.565 mol B. 3.13 mol C. 6.26 mol D. 12.52 mol

OpenStudy (anonymous):

I couldnt figure it out. It's confusing me. I was out of school for 4 days due to the stomach virus.

OpenStudy (sleepyjess):

Welcome to OpenStudy! This should probably go in the \(\href{http:///openstudy.com/study#/groups/Chemistry}{\sf Chemistry}\) section :)

OpenStudy (anonymous):

Theres no one there to help. And i really need it done.

OpenStudy (sleepyjess):

@asnaseer ?

OpenStudy (anonymous):

Mol ratio is 1:1 From looking at the balanced equation, for every 1 mol of CO2, we produce 1 mol Li2CO3. Stoichiometry \[\frac{ 3.13 mol CO2 }{ } \times \frac{ 1 mol Li2CO3 }{1 mol CO2 } = 3.13 mol Li2CO3\] I suggest you use stoichiometry every time you these problems, as it gets complicated later if you need to get mols to grams to molarity and so forth.